the sample mean is computed from a random sample of size 25 drama from a population having SD find i)90% ii) 99% confidence interval of the population mean


1
Expert's answer
2021-05-07T14:16:17-0400

Sample size "n = 25"

Sample mean "= \bar x"

Sample standard deviation "= s"


"CI = \bar x \pm z \dfrac{s}{\sqrt{n}}"


a.) "CI = \bar x \pm 1.645 \dfrac{s}{\sqrt{n}}" (z score at 90% = 1.645 )


b.) "CI = \bar x \pm 1.96 \dfrac{s}{\sqrt{n}}" (z score at 95% = 1.96 )